# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v3

@@ Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly used for odometry and SLAM (Simultaneous Location and Mapping). However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing registration algorithms, especially those originally intended for more accurate sensors such as LiDAR. Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent probability distribution function for registration. Moreover, we propose tackling the problem of radar noise by optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function. Finally, following existing practice we implement an Extended Kalman Filter-based Radar-Inertial Odometry pipeline in order to evaluate the effectiveness of our system. Experiments using publicly available 4D radar datasets show that our Gaussian approach is comparable to existing registration algorithms, outperforming them in several sequences.
@@ I Introduction
Figure 1: Gaussian model of an example point cloud, containing 150 Gaussians. The surface of each spheroid represents the $2\sigma$ front of each Gaussian (i.e. Mahalanobis distance equal to two).
Figure 2: Overview of the Radar-Inertial Odometry pipeline including our proposed Gaussian modeling and scan matching processes
We introduce a new method for modeling and registering 4D radar point clouds using 3D Gaussians. Following the success of 3D Gaussian Splatting (3DGS) for vision [6], we propose the use of a collection of 3D Gaussians for modeling the geometry of a scene (see Fig. 1). As in [6] (and in contrast to NDT [1]), we simultaneously optimize all parameters of the model, including the position, scaling, rotation and correspondences between points and Gaussians. This approach is entirely featureless, and does not require any priors other than an initial set of centroids. Moreover, we propose a registration step that considers multiple simultaneous hypotheses in order to add further robustness against radar noise. Finally, we integrate our algorithms into a Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) (see Fig. 2), which we then use to evaluate our Gaussian methods against several established registration algorithms on multiple datasets and radar types.
@@ II Related work
The Normal Distributions Transform (NDT) algorithm [1] subdivides a point cloud using a voxel grid and computes normal distributions for each cell that contains points. Scan matching is performed by maximizing the resulting probability distribution function with respect to an incoming scan. An important limitation of this algorithm is that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment. Another problem lies in that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization. NDT has been used with radar scans too [13, 14]. Voxelized GICP (VGICP) [15] is a derivative that combines NDT’s voxel-based representation with the distribution-to-distribution based approach of GICP for registration. G3Reg [16] is another derivative of NDT, which performs classification on each voxel (plane/cluster/point) in order to merge plane cells together, and extract semantic information to further aid scan matching. In all these algorithms, each individual Gaussian is modeled separately using their associated subcloud extracted from the segmentation. However, we are interested in approaches that jointly optimize all Gaussians, looking at the entire point cloud.
@@ III Gaussian modeling
Given a point cloud $\mathcal{P}$ contaning $M$ points $\{\mathbf{p}_{i}|i=0,1,\ldots,M\}$ with $\mathbf{p}_{i}\in\mathbb{R}^{3}$, our goal is creating a summarized representation of its geometry for use with downstream tasks such as mapping or localization. We model the geometry of a given radar scene using normal distributions, as in [1, 6, 15, 16]. Figure 1 shows an example of a Gaussian model.
@@ III-A Parametrization
  • Scaling vector $\mathbf{s}_{j}\in\mathbb{R}^{3}$, indicating the natural logarithms of the square roots of the eigenvalues of the covariance matrix. We impose a minimum size restriction so that Gaussians do not become infinitely small (and Mahalanobis distances infinitely large).
  • Rotation quaternion $\mathbf{q}_{j}\in\mathbb{H}$, compactly encoding the eigenvectors of the covariance matrix (see below); where $\mathbb{H}$ is the field of quaternions. In order to ensure that the quaternions remain unitary during the optimization, we normalize them prior to usage.
Unlike in Gaussian Mixture Models (GMMs), we do not define $\pi_{j}$ weight parameters for the prior probability of a point being generated by a given Gaussian. This is because point cloud data does not encode a complete population of points – it merely represents a sample of said population, biased by factors such as the physical geometric properties of the sensor. For instance, sampling density is inversely proportional to distance from the sensor. This means we cannot model these prior probabilities, and as such we only consider posterior probabilities after a point is assigned to a Gaussian (through other non-probabilistic means, such as measuring distances).
The scaling and rotation parameters succinctly describe the covariance matrix of a normal distribution [26, 27]. To see why this is possible, we first derive a rotation matrix $\mathbf{R}_{j}=\mathbf{R}\{\mathbf{q}_{j}\}$, and a scaling matrix $\mathbf{S}_{j}=\mathrm{diag}(\exp_{\circ}(\mathbf{s}_{j}))$, where $\exp_{\circ}(\mathbf{x})$ is element-wise natural exponentiation and $\mathbf{R}\{\mathbf{q}\}$ is the quaternion-to-rotation-matrix conversion operator. We then define a transformation matrix $\mathbf{M}_{j}=\mathbf{R}_{j}\mathbf{S}_{j}$ that maps a mean-centered point into the Gaussian’s correlation space, and finally derive the covariance matrix $\boldsymbol{\Sigma}_{j}=\mathbf{M}_{j}\mathbf{M}_{j}^{T}$ of the distribution.
@@ III-B Initialization
First, it is necessary to determine $N$, that is, the number of 3D Gaussians in the model. We propose choosing $N$ proportionally based on the number of points, so that richer scans can be represented with more Gaussians. The model $\mathcal{G}$ is then initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\mathbf{q}_{j}$ set to the unit quaternion. As for $\boldsymbol{\mu}_{j}$, we perform Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
@@ III-C Optimization
We define the following loss function affecting a given $\theta_{j}$, derived from the log PDF (Probability Density Function) of the multivariate normal distribution:
@@ IV Gaussian multi-hypothesis scan matching
Given a Gaussian model $\mathcal{G}$ and a point cloud $\mathcal{P}$, we define a registration process that estimates the pose of the robot $\xi$ with respect to $\mathcal{G}$. The algorithm is based on the simultaneous optimization of multiple hypotheses centered around the current predicted position of the robot, in order to increase the robustness against local optima.
@@ IV-C Keyframing
  • For the translational component: $\mathbf{t}||_{2}\geq d_{\max}$, where $d_{\max}$ is the maximum distance threshold.
  • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation angle threshold.
@@ V-C EKF update and error state reset
Finally, in order to prevent especially noisy observations from causing instability in the filter, we perform Mahalanobis outlier rejection, using a threshold calculated from a given percentile of the $\chi^{2}$ distribution, like [29].
@@ V-D Radar egovelocity observation
Following previous works, we directly observe the robot’s velocity by leveraging the Doppler effect in measurements returned by the radar. We use the RANSAC-LSQ algorithm [29] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot, including a covariance matrix. The filtered point cloud is also fed to the Gaussian modeling and scan matching algorithm in the following subsection. The observation model used to perform the EKF update is as follows:
@@ V-E Scan matching observation
The Gaussian modeling and scan matching algorithms allow us to observe robot poses relative to previous keyframes. In other words, given a keyframe pose $\xi^{k}$ we can observe the change in translation and rotation that has occurred since said keyframe. We initialize the pose particle swarm with a distribution $\mathbf{P}\thicksim\mathcal{N}(\xi^{x},\boldsymbol{\Sigma})$ where $\xi^{x}$ is the relative robot pose formed from the current nominal state of the EKF: $\xi^{x}=\{\mathbf{p},\mathbf{q}\}\ominus\xi^{k}$, and $\boldsymbol{\Sigma}$ is a covariance matrix indicating the desired dispersion of the particles for each of the 6 degrees of freedom (X/Y/Z/roll/pitch/yaw). Afterwards, we apply the scan matching algorithm and obtain an optimized observation pose $\xi^{y}$, from which a residual pose can be obtained: $\xi^{r}=\xi^{y}\ominus\xi^{x}=\{\delta\mathbf{p},\delta\mathbf{q}\}$. Since the attitude error quaternion $\delta\mathbf{q}$ can be approximated as $\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right)\approx[1\,\frac{1}{2}\delta\boldsymbol{\theta}^{T}]^{T}$ for very small errors, the attitude error vector can be approximated as $\delta\boldsymbol{\theta}\approx\frac{2}{\delta q_{w}}\delta\mathbf{q}_{v}$. This allows us to construct the residual vector $\mathbf{r}=[\delta\mathbf{p}^{T}\,\delta\boldsymbol{\theta}^{T}]^{T}$ to be fed to the EKF update equation – note that this vector contains X/Y/Z/roll/pitch/yaw components (in this order). The non-zero subblocks of the observation matrix $\mathbf{H}$ can be defined as such: $\mathbf{H}_{\delta p\,p}=\mathbf{H}_{\delta\theta\,\delta\theta}=\mathbf{C}_{w}^{k}$, where $k$ corresponds to the keyframe’s reference frame. We are currently using a fixed observation covariance matrix $\mathbf{R}$, although we believe it should be possible to leverage the optimized particle swarm returned by the scan matching algorithm to estimate this matrix.
Table I: Results on NTU4DRadLM Dataset (Oculii Eagle). Best results in bold, second best results underlined.
Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching, Abl.: Ablation
cp nyl loop2 loop3
Usage Time $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
IMU EV SM (ms) (%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
@@
Table II: Results on Snail-Radar Dataset (Oculii Eagle). Best results in bold, second best results underlined.
st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
@@
Table III: Results on Snail-Radar Dataset (ARS548). Best results in bold, second best results underlined.
st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
@@
Furthermore, we constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor. For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms. In order to ignore the unreliable degrees of freedom, we transform the residual vector, observation matrix and covariance matrix by removing the rows and/or columns corresponding to Z/roll/pitch prior to applying the EKF update.
@@ VI Experimental results
Besides our Gaussian methods, we evaluate four ablations within our RIO framework: one without multi-hypothesis registration, and three using common established scan matching algorithms (NDT, GICP and VGICP) directly with radar point clouds. We use the current nominal state of the EKF directly as the initial estimate for single-hypothesis scan matching. The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [33] as the underlying implementation. In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian. In both cases we sample a swarm of 8 particles during scan matching.

### v4

@@ Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly used for odometry and SLAM (Simultaneous Location and Mapping). However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing registration algorithms, especially those originally intended for more accurate sensors such as LiDAR. Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent Probability Density Function (PDF) for registration. Moreover, we propose tackling the problem of radar noise entirely within the scan matching process by optimizing multiple registration hypotheses for better protection against local optima of the PDF. Finally, following existing practice we implement an Extended Kalman Filter-based Radar-Inertial Odometry pipeline in order to evaluate the effectiveness of our system. Experiments using publicly available 4D radar datasets show that our Gaussian approach is comparable to existing registration algorithms, outperforming them in several sequences.
@@ I Introduction
Figure 1: Overview of the Radar-Inertial Odometry pipeline including our proposed Gaussian modeling and scan matching processes. Bold font indicates input/output, the highlighted area indicates our contribution
We introduce a new method for modeling and registering 4D radar point clouds using 3D Gaussians. Following the success of 3D Gaussian Splatting (3DGS) for vision [6], we propose the use of a collection of 3D Gaussians for modeling the geometry of a scene. As in [6] (and in contrast to NDT [1]), we simultaneously optimize all parameters of the model, including the position, scaling, rotation and correspondences between points and Gaussians. This approach is entirely featureless, and does not require any priors other than an initial set of centroids. Moreover, we propose a registration step that considers multiple simultaneous hypotheses in order to add further robustness against radar noise. Finally, we integrate our algorithms into a Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) (see Fig. 1), which we then use to evaluate our Gaussian methods against several established registration algorithms on multiple datasets and radar types.
@@ II Related work
The Normal Distributions Transform (NDT) algorithm [1] subdivides a point cloud using a voxel grid and computes normal distributions for each cell that contains points. Scan matching is performed by maximizing the resulting Probability Density Function (PDF) with respect to an incoming scan. An important limitation of this algorithm is that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment. Another problem lies in that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization. NDT has been used with radar scans too [13, 14]. Voxelized GICP (VGICP) [15] is a derivative that combines NDT’s voxel-based representation with the distribution-to-distribution based approach of GICP for registration. G3Reg [16] is another derivative of NDT, which performs classification on each voxel (plane/cluster/point) in order to merge plane cells together, and extract semantic information to further aid scan matching. In all these algorithms, each individual Gaussian is modeled separately using their associated subcloud extracted from the segmentation. However, we are interested in approaches that jointly optimize all Gaussians, looking at the entire point cloud.
@@ III Gaussian modeling
Given a point cloud $\mathcal{P}$ containing $M$ points $\{\mathbf{p}_{i}|i=0,1,\ldots,M\}$ with $\mathbf{p}_{i}\in\mathbb{R}^{3}$, our goal is creating a summarized representation of its geometry for use with downstream tasks such as mapping or localization. We model the geometry of a given radar scene using normal distributions, as in [1, 6, 15, 16].
@@ III-A Parametrization
  • Scaling vector $\mathbf{s}_{j}\in\mathbb{R}^{3}$, indicating the standard deviations in each dimension. We impose a minimum size restriction so that Gaussians do not become infinitely small (and Mahalanobis distances infinitely large).
  • Rotation quaternion $\mathbf{q}_{j}\in\mathbb{H}$, compactly encoding the orientation of the Gaussian. In order to ensure that the quaternions remain unitary during the optimization, we normalize them prior to usage.
Unlike Gaussian Mixture Models (GMMs), we do not define $\pi_{j}$ weight parameters for the prior probability of a point being generated by a given Gaussian. This is because point cloud data does not encode a complete population of points – it merely represents a sample of said population, biased by factors such as the physical geometric properties of the sensor. For instance, sampling density is inversely proportional to distance from the sensor. This means we cannot model these prior probabilities, and as such we only consider posterior probabilities after a point is assigned to a Gaussian (using e.g. the shortest Euclidean distance).
The scaling and rotation parameters succinctly describe the covariance matrix of a normal distribution [26, 27]. We first derive a rotation matrix $\mathbf{R}_{j}=\mathbf{R}\{\mathbf{q}_{j}\}$, and a scaling matrix $\mathbf{S}_{j}=\mathrm{diag}(\exp_{\circ}(\mathbf{s}_{j}))$, where $\exp_{\circ}(\mathbf{x})$ is element-wise natural exponentiation and $\mathbf{R}\{\mathbf{q}\}$ is the rotation matrix derived from a given quaternion. We then define a matrix $\mathbf{M}_{j}=\mathbf{R}_{j}\mathbf{S}_{j}$ that maps a mean-centered point into the Gaussian’s correlation space, and finally derive the covariance matrix $\boldsymbol{\Sigma}_{j}=\mathbf{M}_{j}\mathbf{M}_{j}^{T}$ of the distribution.
@@ III-B Initialization
First, it is necessary to determine $N$, that is, the number of 3D Gaussians in the model. We propose choosing $N$ proportionally based on the number of points, so that richer scans can be represented with more Gaussians. The proportion should also consider the sparsity of the point clouds generated by the radar used. The model $\mathcal{G}$ is then initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\mathbf{q}_{j}$ set to the unit quaternion. As for $\boldsymbol{\mu}_{j}$, we perform Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
@@ III-C Optimization
We define the following loss function for each $\theta_{j}$, derived from the log PDF of the multivariate normal distribution:
@@ IV Gaussian multi-hypothesis scan matching
Given a Gaussian model $\mathcal{G}$ and a point cloud $\mathcal{P}$, we define a registration process that estimates the pose of the robot $\xi$ with respect to $\mathcal{G}$. The algorithm is based on the simultaneous optimization of multiple hypotheses centered around the current predicted position of the robot, in order to increase the robustness against local optima. At the end of the optimization, the hypotheses are judged against a scoring function, and the best hypothesis is selected as the output of the scan matching algorithm while the other hypotheses are discarded. Note that this is unlike a classical particle filter, which does persist the hypotheses throughout time.
@@ IV-C Keyframing
  • For the translational component: $\mathbf{t}||_{2}\geq{t}_{\max}$, where ${t}_{\max}$ is the maximum translation threshold.
  • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation threshold.
@@ V-C EKF update and error state reset
Finally, in order to prevent especially noisy observations from causing instability in the filter, we follow [29] and perform Mahalanobis outlier rejection, using a threshold calculated from a given percentile of the $\chi^{2}$ distribution.
@@ V-D Radar egovelocity observation
Following previous works, we directly observe the robot’s velocity by leveraging the Doppler effect in measurements returned by the radar. We use the RANSAC-LSQ algorithm [29] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot, including a covariance matrix. The filtered point cloud is then fed to the Gaussian modeling and scan matching algorithm in the following subsection. The observation model used to perform the EKF update is as follows:
@@ V-E Scan matching observation
The Gaussian modeling and scan matching algorithms allow us to observe robot poses relative to previous keyframes. In other words, given a keyframe pose $\xi^{k}$ we can observe the change in translation and rotation that has occurred since said keyframe. We randomly initialize the pose particle swarm with a distribution $\mathbf{P}\thicksim\mathcal{N}(\xi^{x},\boldsymbol{\Sigma})$ where $\xi^{x}$ is the relative robot pose formed from the current nominal state of the EKF: $\xi^{x}=\{\mathbf{p},\mathbf{q}\}\ominus\xi^{k}$, and $\boldsymbol{\Sigma}$ is a covariance matrix indicating the desired dispersion of the particles for each of the 6 degrees of freedom (X/Y/Z/roll/pitch/yaw). Afterwards, we apply the scan matching algorithm and obtain a single optimized observation pose $\xi^{y}$, from which a residual pose can be obtained: $\xi^{r}=\xi^{y}\ominus\xi^{x}=\{\delta\mathbf{p},\delta\mathbf{q}\}$. Since the attitude error quaternion $\delta\mathbf{q}$ can be approximated as $\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right)\approx[1\,\frac{1}{2}\delta\boldsymbol{\theta}^{T}]^{T}$ for very small errors, the attitude error vector can be approximated as $\delta\boldsymbol{\theta}\approx\frac{2}{\delta q_{w}}\delta\mathbf{q}_{v}$. This allows us to construct the residual vector $\mathbf{r}=[\delta\mathbf{p}^{T}\,\delta\boldsymbol{\theta}^{T}]^{T}$ to be fed to the EKF update equation – note that this vector contains X/Y/Z/roll/pitch/yaw components (in this order). The non-zero subblocks of the observation matrix $\mathbf{H}$ can be defined as such: $\mathbf{H}_{\delta p\,p}=\mathbf{H}_{\delta\theta\,\delta\theta}=\mathbf{C}_{w}^{k}$, where $k$ corresponds to the keyframe’s reference frame. We currently use a fixed observation covariance matrix $\mathbf{R}$, although we believe it is possible to leverage the optimized particle swarm returned by the scan matching algorithm to estimate this matrix.
Table I: Results on NTU4DRadLM Dataset (Oculii Eagle). Best results in bold, second best results underlined. Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching, Abl.: Ablation. Relative translation errors ($t_{\mathrm{rel}}$) are expressed in %, relative rotation errors ($r_{\mathrm{rel}}$) in °/m, and time in ms
Radar/Method Inputs cp nyl loop2 loop3
Oculii Eagle IMU EV SM Time $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
@@
Table II: Results on Snail-Radar Dataset (Oculii Eagle and ARS548). Best results in bold, second best results underlined. Relative translation errors ($t_{\mathrm{rel}}$) are expressed in % and relative rotation errors ($r_{\mathrm{rel}}$) in mrad/m
Radar/Method st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
Oculii Eagle $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
@@
ARS548 $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
@@
Furthermore, we follow [31] and constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor. For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms. In order to ignore the unreliable degrees of freedom, we transform the residual vector, observation matrix and covariance matrix by removing the rows and/or columns corresponding to Z/roll/pitch prior to applying the EKF update.
@@ VI Experimental results
Besides our Gaussian methods, we evaluate four ablations within our RIO framework: one without multi-hypothesis registration, and three using common established scan matching algorithms (NDT, GICP and VGICP) directly with radar point clouds. We use the current nominal state of the EKF directly as the initial estimate for single-hypothesis scan matching. The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [34] as the underlying implementation. In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian. In both cases we sample a swarm of 8 particles during scan matching, which we believe are enough for testing purposes without significantly increasing processing times. Other parameters include $t_{\mathrm{max}}$ = 15m and $\alpha_{\mathrm{max}}$ = 5° for keyframing, $d_{\mathrm{max}}$ = 4 for scan matching, and 0.5m/2° for multi-hypothesis particle dispersion. These parameters were chosen to suit the typical depth range/sparsity relation of a radar.